how do i solve radical63?
Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

sqrt(63)
=sqrt(9*7)
=3*sqrt(7)
.
Ed

Answer by stanbon(75887)

solve radical63
-----------
sqrt[63] = sqrt[9*7] = sqrt(9)*sqrt(7) = 3*sqrt(7)
=============
Cheers,
Stan H.
=============